# Neural Geometry Processing via Spherical Neural Surfaces

### v1

@@
Abstract.
Neural surfaces (e.g., neural map encoding, deep implicits and neural radiance fields) have recently gained popularity because of their generic structure (e.g., multi-layer perceptron) and easy integration with modern learning-based setups. Traditionally, we have a rich toolbox of geometry processing algorithms designed for polygonal meshes to analyze and operate on surface geometry. However, neural representations are typically discretized and converted into a mesh, before applying any geometry processing algorithm. This is unsatisfactory and, as we demonstrate, unnecessary. In this work, we propose a spherical neural surface representation (a spherical parametrization) for genus-0 surfaces and demonstrate how to compute core geometric operators directly on this representation. Namely, we show how to construct the normals and the first and second fundamental forms of the surface, and how to compute the surface gradient, surface divergence and Laplace Beltrami operator on scalar/vector fields defined on the surface. These operators, in turn, enable us to create geometry processing tools that act directly on the neural representations without any unnecessary meshing. We demonstrate illustrative applications in (neural) spectral analysis, heat flow and mean curvature flow, and our method shows robustness to isometric shape variations. We both propose theoretical formulations and validate their numerical estimates. By systematically linking neural surface representations with classical geometry processing algorithms, we believe this work can become a key ingredient in enabling neural geometry processing.
neural representation, continuous surface representation, spectral analysis, meshfree representation, neural geometry processing
submissionid: 0461journal: TOGccs: Computing methodologies Shape analysisccs: Computing methodologies Mesh geometry modelsccs: Computing methodologies Shape representationsccs: Computing methodologies Machine learning approaches
1. Introduction
The emergence of neural networks to represent surfaces has rapidly gained popularity as a generic surface representation. Many variants (Park et al., 2019; Mescheder et al., 2018; Groueix et al., 2018; Qi et al., 2016) have been proposed to directly represent a single surface, as an overfitted network, or a collection of surfaces, as encoder-decoder networks. The ‘simplicity’ of the approach lies in the generic network architectures (e.g., MLPs, UNets) being universal function approximators. Such deep representations can either be overfitted to given meshes (Davies et al., 2020; Morreale et al., 2021) or directly optimized from a set of posed images (i.e., images with camera information) via an optimization, as popularized by the Nerf formulation (Mildenhall et al., 2021). Although there are some isolated and specialized attempts (Novello et al., 2022; Chetan et al., 2023; Yang et al., 2021), no established set of operators directly works on such neural objects, encoding an explicit surface, to facilitate easy integration with existing geometry processing tasks.
@@ 1. Introduction
Meshes, however, are discretized surface representations, and accurately carrying over definitions from continuous to discretized surfaces is surprisingly difficult. For example, the area of discrete differential geometry (DDG) (Desbrun et al., 2006) was entirely invented to meaningfully carry over continuous differential geometry concepts to the discrete setting. However, many commonly used estimates vary with the mesh quality (i.e., distribution of vertices and their connections), even when encoding the same underlying surface. In the case of the Laplace Beltrami operator, the ‘no free lunch’ result (Wardetzky et al., 2007) tells us that we cannot hope for any particular discretization of the Laplace Beltrami operator to simultaneously satisfy all of the properties of its continuous counterpart. In practice, this means that the most suitable choice of discretization (see (Bunge et al., 2023) for a recent survey) may change depending on the specifics of the target application.
@@
In the absence of native geometric operators for neural representations, the common workaround for applying a geometry processing algorithm to neural representations is discretizing these representations into an explicit mesh (e.g., using Marching cubes) and then relying on traditional mesh-based operators. Unfortunately, this approach inherits the disadvantages of having an explicit mesh representation, as listed above. A similar motivation prompted researchers to pursue a grid-free approach for volumetric analysis using Monte Carlo estimation for geometry processing (Sawhney and Crane, 2020) or neural field convolution (Nsampi et al., 2023) to perform general continuous convolutions with continuous signals such as (overfitted) neural fields.
@@
We avoid discretization altogether. First, we create a smooth and differentiable representation — spherical neural surfaces — where we encode individual genus-0 surfaces using dedicated neural networks. Then, we exploit autograd to compute surface normals, tangent planes, and Fundamental Forms directly with continuous differential geometry. Knowing the First Fundamental Form allows us to compute attributes of the parametrization, such as local area-distortion. Knowing the Second Fundamental form allows us to compute surface attributes such as mean curvature, Gauss curvature, and principal curvature directions — without any unnecessary discretization error. Essentially, encoding surfaces as seamless neural maps, almost trivially, allows us to access definitions from differential geometry.
@@
Other estimates require more thought. We use the computed surface normals and mean curvatures to apply the continuous Laplace Beltrami operator to any differentiable scalar/vector field. Additionally, to enable spectral analysis on the resultant Laplace Beltrami operator, we propose a novel optimization scheme and a neural representation of scalar fields that allows us to extract the lowest $k$ spectral modes, using the variational formulation of the Dirichlet eigenvalue problem. We use Monte Carlo estimates to approximate the continuous inner products and integration of functions on the (neural) surface. Our framework of spherical neural surfaces makes optimizing many expressions involving surface gradients and surface integrals of smooth functions possible. We hope that our setup opens the door to solving other variational problems posed on smooth (neural) surfaces.
@@
We evaluate our proposed framework on standard geometry processing tasks. In the context of mesh input, we assess the robustness of our estimates to different meshing of the same underlying surfaces (Figure 11) and isometrically related shapes (Figure 10). We compare the results against ground truth values when they are computable (e.g., analytical functions - see Figure 3 and Figure 4) and contrast them against alternatives, where applicable (Figure 11).
@@
Our main contribution is introducing a novel representation and associated operators to enable neural geometry processing.
We:
@@
    introduce spherical neural surfaces as a representation to encode genus-0 surfaces seamlessly as overfitted neural networks;
@@
    demonstrate the robustness and accuracy of the setup on different forms of input.
2. Related Works
Neural representations
@@ 2. Related Works
Implicit neural representations, such as signed distance functions (SDFs) and occupancy networks, have recently been popularized to model 3D shapes. Notably, DeepSDF (Park et al., 2019) leverages a neural network to represent the SDF of a shape, enabling continuous and differentiable shape representations that can be used for shape completion and reconstruction tasks; (Davies et al., 2020) use neural networks to overfit to individual SDFs; Occupancy Networks (Mescheder et al., 2018) learn a continuous function to represent the occupancy status of points in space, providing a powerful tool for 3D shape modeling and reconstruction from sparse input data. Explicit neural representations, on the other hand, use neural networks to directly predict structured 3D data such as meshes or point clouds. Mesh R-CNN (Gkioxari et al., 2019) extends Mask-RCNN to predict 3D meshes from images by utilizing a voxel-based representation followed by mesh refinement with a graph convolution network operating over the mesh’s vertices and edges. In the context of point clouds, PointNet (Qi et al., 2016) and its variants are widely used as backbones for shape understanding tasks like shape classification and segmentation. Hybrid representations combine the strengths of both implicit and explicit methods. For example, Pixel2Mesh (Wang et al., 2018) generates 3D meshes from images by progressively deforming a template mesh using a graph convolutional network, integrating the detailed geometric structure typical of explicit methods with the continuous nature of implicit representations. More relevant to our work is the AtlasNet (Groueix et al., 2018) that models surfaces as a collection of parametric patches, balancing flexibility and precision in shape representation, and Neural Surface Maps (Morreale et al., 2021), which are overfitted to a flattened disc parametrization of surfaces to enable surface-to-surface mapping.
Estimating differential quantities
Traditionally, researchers have investigated how to carry over differential geometry concepts (do Carmo, 1976) to surface meshes, where differential geometry does not directly apply because mesh faces are flat, with all the ‘curvature’ being at sharp face intersections. Taubin (1995) introduce several signal processing operators on meshes. Meyer et al. (2002) used averaging Voronoi cells and the mixed Finite-Element/Finite-Volume method; (Cazals and Pouget, 2005) used osculating jets; while (de Goes et al., 2020) used discrete differential geometry (Desbrun et al., 2006) to compute gradients, curvatures, and Laplacians on meshes. Recently, researchers have used learning-based approaches to ‘learn’ differential quantities like normals and curvatures on point clouds (Guerrero et al., 2018; Ben-Shabat et al., 2019; Pistilli et al., 2020).
More related to ours is the work by Novello et al. (2022), who represent surfaces via implicit functions and analytically compute differential quantities such as normals and curvatures. With a similar motivation, Chetan et al. (2023) fit local polynomial patches to obtain more accurate derivatives from pre-trained hybrid neural fields; they also use these as higher-order constraints in the neural implicit fitting stage. A similar idea was explored by Bednarik et al. (2020), as a regularizer for the AtlasNet setup, by using implicit differential surface properties during training. The work by Yang and colleagues (2021) is closest to ours in motivation - they also mainly focus on normal and curvature estimates but they additionally demonstrate how to perform feature smoothing and exaggeration using the local differential quantities in the target loss functions and thus driving the modification of the underlying neural fields. However, we go beyond normal and curvature estimates and focus on operators such as surface gradient, surface divergence and Laplace Beltrami operators for processing scalar and vector fields defined on the surfaces.
Laplace Beltrami operators on meshes
@@
The Laplace-Beltrami operator (LBO) is an indispensable tool in spectral geometry processing, as it can be used analyze and manipulate the intrinsic properties of shapes. Spectral mesh processing (Lévy and Zhang, 2010) leverages the eigenvalues and eigenfunctions of the LBO for tasks such as mesh smoothing and shape segmentation. Many subsequent efforts have demonstrated the use of LBO towards shape analysis and understanding (e.g., shapeDNA (Reuter et al., 2006), global point signatures (Rustamov, 2007), heat kernel signature (Sun et al., 2009)), eventually culminating into the functional map framework (Ovsjanikov et al., 2012). An earlier signal processing framework (Taubin, 1995) has been revisited with the LBO operator (Lévy and Zhang, 2010) for multi-scale processing applications (e.g., smoothing and feature enhancement). Although the Laplace-Beltrami operator on triangular meshes or, more generally, on polyhedral meshes, can be computed, these operators depend on the underlying discretization of the surface (i.e., placement of vertices and their connectivity). See (Bunge et al., 2023) for a discussion. In the case of triangular meshes, the commonly used discretizations are the Uniform LBO and the cotan LBO. In Section 6, we compare our neural LBO with the cotan discretization and discuss their relative merits.
3. Spherical Neural Surfaces
@@
We introduce a new neural representation called a ‘Spherical Neural Surface’ (SNS) to represent continuous surfaces. An SNS is a Multi-Layer Perceptron (MLP) $S_{\theta}:\,\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$, which we train so that the restriction to the unit sphere ($\mathbb{S}^{2}\subset\mathbb{R}^{3}$) is a continuous parametrization of a given surface. In terms of notation, we will also use $S_{\theta}$ or $S$ as shorthand for the set $S_{\theta}(\mathbb{S}^{2})$. In our current work, we primarily focus on creating SNSs from input triangle-meshes and analytically-defined surfaces, but there is potential to create SNSs from other representations such as pointclouds, SDFs, etc. Figure 12 shows initial results for generating an SNS from a neural SDF.
3.1. Overfitting an SNS to a Triangle-Mesh
@@
Given a genus-0 manifold triangle-mesh, we first find an injective embedding of the mesh vertices onto the unit sphere, using a spherical embedding algorithm (Schmidt et al., 2023; Praun and Hoppe, 2003). We extend the embedding to a continuous embedding by projecting points on the sphere to the discrete sphere, and employing barycentric coordinates so that every point $p_{i}$ on the sphere corresponds to a unique point $q_{i}$ on the target surface. We overfit the network $S_{\theta}$ to this parametrization by minimizing the MSE between the ground truth and predicted surface positions, as:
(1) $\mathcal{L}_{MSE}:=\frac{1}{N}\sum\limits_{i=1}^{N}\left\|S_{\theta}(p_{i})-q_{i}\right\|^{2}.$
@@
(2) $\mathcal{L}_{\textit{normal}}:=\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{n}_{S_{\theta}}({p_{i}})-\mathbf{n}_{\textit{mesh}}({p_{i}})\right\|^{2}=\frac{2}{N}\sum_{i=1}^{N}(1-\cos{\alpha_{i}(p_{i})}).$
In this expression, $\mathbf{n}_{\textit{mesh}}({p_{i}})$ is the (outwards) unit-normal on the mesh, and $\mathbf{n}_{S_{\theta}}({p_{i}})$ is the corresponding (outwards) unit-normal of the smooth parametrization at $p_{i}$, which is derived analytically from the Jacobian of $S_{\theta}$ at $p_{i}$ (see Section 3.2); the angle $\alpha(p_{i})$ is the angle between $\mathbf{n}_{\textit{mesh}}({p_{i}})$ and $\mathbf{n}_{S_{\theta}}({p_{i}})$. See Figure 2 for an example.
Figure 2. Spherical Neural Surface. Given an input surface — a mesh in this case — we progressively overfit an MLP to represent the surface as $S_{\theta}$. The loss term is simply the MSE between the ground truth and predicted surface positions, plus a (scaled) normals regularization term (Equation 2).
3.2. Computing Differential Quantities using SNS
One of the advantages of Spherical Neural Surfaces as a geometry representation is that it is extremely natural to compute many important quantities from continuous differential geometry - without any need for approximation or discretization. This is thanks to the automatic differentiation functionality built into modern machine learning setups and algebraically tracking some changes of variables.
For example, to compute the outwards-pointing normal $\mathbf{n}_{S}$, we simply compute the $3\times 3$ Jacobian of $S$$\mathbf{J}({S})$ — and then turn this into a Jacobian for a local 2D parametrization by composing it with a $3\times 2$ orthonormal matrix ($R$):
(3) $||\mathbf{J}^{\textit{local}}_{S}=\mathbf{J}({S})\,R=\begin{pmatrix}\textbar&\textbar\\ S_{u}&S_{v}\\ \textbar&\textbar\\ \end{pmatrix}.$
The columns of $\mathbf{J}^{\textit{local}}_{S}$ are the partial derivatives $S_{u}$ and $S_{v}$. The matrix $R$ defines orthogonal unit vectors on the tangent plane of the sphere at each point, and the functions $u$ and $v$ are the corresponding orthonormal coordinates on the tangent plane of the sphere. Then, the normalized cross product of $S_{u}$ and $S_{v}$ is the outward-pointing unit normal to the surface:
(4) $\mathbf{n}_{S}=\frac{S_{u}\times S_{v}}{\left\|S_{u}\times S_{v}\right\|}.$

### v2

@@
Abstract
Neural surfaces (e.g., neural map encoding, deep implicits and neural radiance fields) have recently gained popularity because of their generic structure (e.g., multi-layer perceptron) and easy integration with modern learning-based setups. Traditionally, we have a rich toolbox of geometry processing algorithms designed for polygonal meshes to analyze and operate on surface geometry. In the absence of an analogous toolbox, neural representations are typically discretized and converted into a mesh, before applying any geometry processing algorithm. This is unsatisfactory and, as we demonstrate, unnecessary. In this work, we propose a spherical neural surface representation for genus-0 surfaces and demonstrate how to compute core geometric operators directly on this representation. Namely, we estimate surface normals and first and second fundamental forms of the surface, as well as compute surface gradient, surface divergence and Laplace Beltrami operator on scalar/vector fields defined on the surface. Our representation is fully seamless, overcoming a key limitation of similar explicit representations such as Neural Surface Maps [MAKM21]. These operators, in turn, enable geometry processing directly on the neural representations without any unnecessary meshing. We demonstrate illustrative applications in (neural) spectral analysis, heat flow and mean curvature flow, and evaluate robustness to isometric shape variations. We propose theoretical formulations and validate their numerical estimates, against analytical estimates, mesh-based baselines, and neural alternatives, where available. By systematically linking neural surface representations with classical geometry processing algorithms, we believe that this work can become a key ingredient in enabling neural geometry processing. Code will be released upon acceptance, accessible from the project webpage.
1 Introduction
The use of neural networks to represent surfaces has gained rapid popularity as a generic surface representation. Many variants of this formulation [PFS19, MON18, GFK18, QSMG] have been proposed, such as directly representing a single surface as an overfitted network, or a collection of surfaces, as encoder-decoder networks. The ‘simplicity’ of the approach lies in the network architectures (e.g., MLPs, UNets) being universal function approximators. Such deep representations can either be overfitted to given meshes [DNJ20, MAKM21] or directly optimized from a set of posed images (i.e., images with camera information), as popularized by the NeRF formulation [MST21]. Although there are some isolated and specialized attempts [NSS22, CYW23, YBHK21], no established set of operators directly works on a neural object, encoding an explicit surface, to facilitate easy integration with existing geometry processing tasks.
@@ 1 Introduction
Meshes, however, are discretized surface representations, and accurately carrying over definitions from continuous to discretized surfaces is surprisingly difficult. For example, the area of discrete differential geometry (DDG) [DPSS06] was entirely invented to meaningfully carry over continuous differential geometry concepts to the discrete setting. However, many commonly used estimates vary with the mesh quality (i.e., distribution of vertices and their connections), even when encoding the same underlying surface. In the case of the Laplace-Beltrami operator, the ‘no free lunch’ result [WMKG07] tells us that we cannot hope for any particular discretization of the Laplace-Beltrami operator to simultaneously satisfy all of the properties of its continuous counterpart. In practice, the most suitable choice of discretization, such as triangular or, more generally, polygonal meshes, can change depending on the specifics of the target application (see [BAB23] for a recent survey).
@@
In the absence of native geometric operators for neural representations, the common workaround for applying a geometry processing algorithm to neural representations is discretizing these representations into an explicit mesh (e.g., using Marching cubes) and then relying on traditional mesh-based operators. Unfortunately, this approach inherits the disadvantages of having an explicit mesh representation, as discussed above. A similar motivation prompted researchers to pursue a grid-free approach for volumetric analysis using Monte Carlo estimation for geometry processing [SC20] or neural field convolution [NDS23] to perform general continuous convolutions with continuous signals such as neural fields.
@@
We propose a different approach to discretization-free geometry processing. First, we create a smooth and differentiable representation — spherical neural surfaces (SNS) — where we encode individual genus-0 surfaces using dedicated neural networks. Having a differentiable encoding allows one to directly compute surface normals, tangent planes, and Fundamental Forms directly with continuous differential geometry. The First Fundamental Form allows us to compute attributes of the parameterization, such as local area distortion. The Second Fundamental form allows us to compute surface attributes such as mean curvature, Gauss curvature, and principal curvature directions. Essentially, encoding surfaces as seamless neural maps unlocks definitions from differential geometry.
@@
Other estimates require more thought. We implement two methods to apply the continuous Laplace Beltrami operator to any differentiable scalar or vector field (using Equations 13 and 16). Additionally, to enable spectral analysis on the resultant Laplace Beltrami operator, we propose a novel optimization scheme via a neural representation of scalar fields that allows us to extract the lowest $k$ spectral modes, using the variational formulation of the Dirichlet eigenvalue problem. We use Monte Carlo estimates to approximate the continuous inner products and integration of functions on the (neural) surface. Our framework of spherical neural surfaces makes optimizing many expressions involving surface gradients and surface integrals of smooth functions possible. We hope that our setup opens the door to solving other variational problems posed on smooth (neural) surfaces.
@@
We evaluate our proposed framework on standard geometry processing tasks. In the context of mesh input, we assess the robustness of our estimates to different meshing of the same underlying surfaces (Figure 10) and isometrically related shapes (Figure 9). We compare the results against ground truth values when they are computable (e.g., analytical functions – see Figure 2 and Figure 3) and contrast them against alternatives. Specifically, we compare against traditional mesh-based methods (e.g., curvature estimates, cotan LBO, – see Figures 8, 10, 11) or neural alternatives (e.g., NGP [YBHK21] and neural implicits [NSS22] – see Figure 6).
@@
Our main contribution is introducing a novel representation and associated operators to enable neural geometry processing. We:
@@
    introduce spherical neural surfaces as a representation to encode genus-0 surfaces seamlessly as overfitted networks;
@@
    demonstrate the robustness and accuracy ours on different forms of input and quantitatively compare against alternatives.
2 Related Works
Neural representations.
@@ 2 Related Works
Implicit neural representations, such as signed distance functions (SDFs) and occupancy networks, have recently been popularized to model 3D shapes. Notably, DeepSDF [PFS19] leverages a neural network to represent the SDF of a shape, enabling continuous and differentiable shape representations that can be used for shape completion and reconstruction tasks; Davies et al. [DNJ20] use neural networks to overfit to individual SDFs; Occupancy Networks [MON18] learn a continuous function to represent the occupancy status of points in space, providing a powerful tool for 3D shape modeling and reconstruction from sparse input data. Explicit neural representations, on the other hand, use neural networks to directly predict structured 3D data such as meshes or point clouds. Mesh R-CNN [GMJ19] extends Mask-RCNN to predict 3D meshes from images using a voxel-based representation followed by mesh refinement with a graph convolution network operating on the meshes’ vertices and edges. In the context of point clouds, PointNet [QSMG] and its variants are widely used as backbones for shape understanding tasks such as shape classification and segmentation. Hybrid representations combine the strengths of both implicit and explicit methods. For example, Pixel2Mesh [WZL18] generates 3D meshes from images by progressively deforming a template mesh using a graph convolutional network, integrating the detailed geometric structure typical of explicit methods with the continuous nature of implicit representations. More relevant to ours is AtlasNet [GFK18] that models surfaces as a collection of parametric patches, balancing flexibility and precision in shape representation, and Neural Surface Maps [MAKM21], which are overfitted to a flattened disc parametrization of surfaces to enable surface-to-surface mapping. Similarly to SNS, Explicit Neural Surfaces [WMVB23] were proposed in the context of 3D reconstruction from images, but the authors did not yet explore the properties of the differentiable surface representation.
Estimating differential quantities.
[…]
Laplace Beltrami operators on meshes.
@@
The Laplace-Beltrami operator (LBO) is an indispensable tool in spectral geometry processing, as it is used to analyze and manipulate the intrinsic properties of shapes. Spectral mesh processing [LZ10] leverages the eigenvalues and eigenfunctions of the LBO for tasks such as mesh smoothing and shape segmentation. Many subsequent efforts have demonstrated the use of LBO towards shape analysis and understanding (e.g., shapeDNA [RWP06], global point signatures [Rus07], heat kernel signature [SOG09]), eventually culminating into the functional map framework [OBCS12]. An earlier signal processing framework [Tau95] has been revisited with the LBO operator [LZ10] for multi-scale processing applications (e.g., smoothing and feature enhancement). Although the Laplace-Beltrami operator on triangular meshes or, more generally, on polyhedral meshes, can be computed, these operators depend on the underlying discretization of the surface (i.e., placement of vertices and their connectivity). See [BAB23] for a discussion. In the case of triangular meshes, the commonly used discretizations are the Uniform LBO and the cotan LBO. In Section 6, we compare our neural LBO with the cotan discretization and discuss their relative merits.
3 Spherical Neural Surfaces
@@
We introduce a novel neural representation for continuous surfaces, called a ‘Spherical Neural Surface’ (SNS). An SNS is a Multi-Layer Perceptron (MLP) $S_{\theta}:\,\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$, trained such that its restriction to the unit sphere ($\mathbb{S}^{2}\subset\mathbb{R}^{3}$) is a continuous parametrization of a given surface. In terms of notation, we use $S_{\theta}$ or $S$ as shorthand for the set $S_{\theta}(\mathbb{S}^{2})$. In our current work, we primarily focus on creating SNSs from input triangle-meshes and analytically-defined surfaces, but there is potential to create SNSs from other representations such as pointclouds, SDFs and NeRFs. For example, Figure 13 shows initial results for generating an SNS from a neural SDF.
Figure 1: Spherical Neural Surface. Given an input surface — a mesh in this case — we progressively overfit an MLP to a spherical mesh-embedding, to represent the surface as $S_{\theta}$. The loss term is simply the MSE between the ground truth and predicted surface positions, plus a (scaled) normals regularization term (Equation 2).
3.1 Overfitting an SNS to a Triangle Mesh
@@
Given a genus-0 manifold triangle mesh, we first find an injective embedding of the mesh vertices onto the unit sphere, using a spherical embedding algorithm [SPK23, PH03]. We extend the embedding to a continuous embedding by employing barycentric coordinates, so that every point $p_{i}$ on a (curved) triangle on the sphere corresponds to a unique point $q_{i}$ on a (flat) triangle on the target surface. We overfit the network $S_{\theta}$ to this parametrization by minimizing the MSE between the ground truth and predicted surface positions:
$\mathcal{L}_{MSE}:=\frac{1}{N}\sum\limits_{i=1}^{N}\left\|S_{\theta}(p_{i})-q_{i}\right\|^{2}.$ (1)
@@
$\mathcal{L}_{\textit{normal}}:=\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{n}_{S_{\theta}}({p_{i}})-\mathbf{n}_{\textit{mesh}}({p_{i}})\right\|^{2}=\frac{2}{N}\sum_{i=1}^{N}(1-\cos{\alpha_{i}(p_{i})}).$ (2)
In this expression, $\mathbf{n}_{\textit{mesh}}({p_{i}})$ is the (outwards) unit-normal on the mesh, and $\mathbf{n}_{S_{\theta}}({p_{i}})$ is the corresponding (outwards) unit-normal of the smooth parametrization at $p_{i}$, which is derived analytically from the Jacobian of $S_{\theta}$ at $p_{i}$ (see Section 3.2); the angle $\alpha(p_{i})$ is the angle between $\mathbf{n}_{\textit{mesh}}({p_{i}})$ and $\mathbf{n}_{S_{\theta}}({p_{i}})$. Figure 1 shows an example of an SNS over training stages.
3.2 Computing Differential Quantities using SNS
One of the advantages of Spherical Neural Surfaces as a geometric representation is that one can compute important quantities from continuous differential geometry - without any need for approximate numerical formulations or discretization. We just need to use the automatic differentiation functionality built into modern machine learning setups, and algebraically track some changes of variables, before applying the formulas from classical differential geometry.
For example, to compute the outwards-pointing normal $\mathbf{n}_{S}$, we first use autodiff to compute the $3\times 3$ Jacobian of $S$,
$\mathbf{J}({S}):=\begin{bmatrix}\frac{\partial{S}}{\partial{x}},&\frac{\partial{S}}{\partial{y}},&\frac{\partial{S}}{\partial{z}}\end{bmatrix},$
and then turn this into a $3\times 2$ Jacobian in local co-ordinates, by composing it with a $3\times 2$ orthonormal matrix ($R$):
$\mathbf{J}^{\textit{local}}_{S}=\mathbf{J}({S})\,R.$ (3)
The matrix $R$ maps vectors in 2D local co-ordinates to tangent vectors on the sphere at a particular point, and the matrix $J(S)$ maps tangent vectors on the sphere to tangent vectors on the surface $S$. If $u$ and $v$ are the corresponding coordinates on the tangent plane of the sphere, then the columns of $\mathbf{J}^{\textit{local}}_{S}$ are equal to the partial derivatives $S_{u}$ and $S_{v}$ respectively:
$||\mathbf{J}^{\textit{local}}_{S}=\begin{pmatrix}\textbar&\textbar\\ S_{u}&S_{v}\\ \textbar&\textbar\\ \end{pmatrix},$ (4)
and these vectors lie in the tangent plane. The normalized cross product of $S_{u}$ and $S_{v}$ is the outward-pointing unit normal:
$\mathbf{n}_{S}=\frac{S_{u}\times S_{v}}{\left\|S_{u}\times S_{v}\right\|}.$ (5)